OpenStudy (anonymous):

What is the value of x? -6x=42

pooja195 (pooja195):

Any ideas on what the first step maybe? :)

OpenStudy (anonymous):

-6 -42?

pooja195 (pooja195):

nope rmember that we are trying to isolate x there is nothing to subtract we would DIVIDE both sides by-6 -6/-6=1 so just x 42/-6=?

pooja195 (pooja195):

*remember

OpenStudy (mathmale):

-6x=42 On the left side we have multiplication: (-6)*x. The inverse of multiplication is division. That is, to "undo" multiplication, we divide. If -6x=42 and we want to solve for x, we "undo" the multiplication of x by (-6) by dividing both sides of this equation by -6.

pooja195 (pooja195):

\[ \frac{ -6x }{ -6 }=\frac{ 42 }{ -6}\]

pooja195 (pooja195):

Try putting that into the equation tool @mathmale

pooja195 (pooja195):

*the thing you were trying to put in

OpenStudy (mathmale):

\[\frac{ -6x }{ -6 }=\frac{ 42 }{ -6 }\]

pooja195 (pooja195):

There you go! :)

OpenStudy (mathmale):

On the left, the -6 in numerator and denominator cancel, leaving us with just x.

OpenStudy (mathmale):

What is the value of\[\frac{ 42 }{ 6 }?.of.\frac{ 42 }{ -6 }?\]

OpenStudy (mathmale):

@tarap0123: ?